# Paloma[Uncaptioned image] : A Benchmark for Evaluating Language Model Fit

### v1

@@ Appendix A Evaluation Data Source Details
Papasavva et al. (2020) collect posts between June 2016 and November 2019 from the Politically Incorrect board (/pol/) of 4chan, a fringe imageboard emphasizing anonymity and ephemerality. Users can post content without registering, with a thread consisting of an image and message followed by a sequence comments. Threads are deleted shortly after they become inactive. As noted previously, 4chan Corpus has toxicity and mysogynist hate comparable to the worst data in Manosphere Corpus and hatespeech above Gab Corpus. We sample this corpus to measure fit to types of discourse and toxicity that can arise from anonymous posting. We concatenate posts in a thread together with post metadata prepended as a header. This source has no marked domains.
Appendix B Reweighting Perplexities
@@
$\text{NLL}_{\text{d,c}}=-\frac{1}{\mathbf{T}(N_{d,c})}\sum_{t\in N_{d,c}}\sum_{i=1}^{\mid t\mid}\text{ln}~{}p(t_{i}|t_{<i})$ (6)
@@
$\alpha_{d,c^{\prime}}=\frac{\mathbf{T}(N_{d,c^{\prime}})}{\sum\limits_{d^{\prime}\in D}\mathbf{T}(N_{d^{\prime},c^{\prime}})}$ (7)
@@
$\text{perplexity}=\text{exp}\left(\sum_{d\in D}\alpha_{d,c^{\prime}}\text{NLL}_{\text{d,c}}\right)$ (8)
Appendix C Baseline Models
The 6 baseline 1B parameter models that we train employ the following architecture: 2048 maximum sequence length, 2048 model dimension, 16 layers, 16 attention heads, RoPE embedding (Su et al., 2021), SwiGLU activation (Shazeer, 2020), mixed precision, non-parametric layer normalization, and sequential model blocks for attention and feed-forward networks. We use EleutherAI’s GPT NeoX tokenizer (Black et al., 2022) but add 3 additional special tokens that are used to mask PII in Dolma. We train to 35k steps ($\sim$150B tokens) with the following LionW optimizer (Chen et al., 2023) configurations: 2.0e-4 peak learning rate, warm-up of 2000 steps, cosine decay to 70k steps ($\sim$300B tokens), 0.1 weight decay, and betas of 0.9 and 0.95. Note that our batch size varies slightly to accommodate two groups of baselines that were run on different hardware. The Dolma and Falcon Refinedweb baselines were run with a batch size of 2112 training instances per step on 24 A100s for 9 days per model. The RedPajama, The Pile, C4, and mC4-en baselines were run with a batch size of 2048 on 64 AMD Instinct MI250X GPUs888amd.com/en/products/server-accelerators/instinct-mi250x for 2 days per model. In each case we save model checkpoints every 5k steps ($\sim$20B tokens).
Appendix D Formatting and Subsampling
@@
We find preliminary evidence that the monotonic decrease in variability with increased evaluation or training data (see §3.2.1) depends on using the non-concatenated inference input format detailed in §3.2.3. In Table 16 we see that the previously observed trends break down when inputs are concatenated. Additionally, the concatenated documents are drawn from 4 random shufflings where the 4 seeds change for each cell. For comparison the bottom of the table shows results when documents are evaluated separately and with the same set of 4 random seeds for all cells. In both input formats documents that are longer than the model context window are split into separate inputs with no overlap.
@@
Appendix E Most and Least Improved Domains
In §4.2 we show that improvement of LM fit when scaling is unequal from domain to domain. Differences in improvement rates can actually indicate several different training dynamics, exemplified in Figure 6. Looking at performance curves over the underlying factor of scale, helps show more specifically what is going on. Examining the domains at the extreme values of improvement rate is one way to surface interesting details of model fit. In Figure 13 we examine performance curves of the most and least improved domains with respect to number of tokens seen, $\Delta_{t}(\sim 20B,\sim 150B)$, and in Figure 14 we examine the most and least improved with respect to number of model parameters, $\Delta_{p}(85M,805M)$ and $\Delta_{p}(805M,6.4B)$.
Figure 13: Perplexity curves for the most and least improved domains over an increase in tokens seen (See §4.2.1). Columns are specific baseline models; rows are specific evaluation sources.
Figure 14: Perplexity curves for the most and least improved domains over an increase in model size (See §4.2.2). Columns are comparisons of specific model sizes. Each row shows first one (left two subplots) and then another (right two subplots) set of evaluation sources.

### v2

@@ Appendix E Evaluation Data Source Details
Papasavva et al. [2020] collect posts between June 2016 and November 2019 from the Politically Incorrect board (/pol/) of 4chan, a fringe imageboard emphasizing anonymity and ephemerality. Users can post content without registering, with a thread consisting of an image and message followed by a sequence comments. Threads are deleted shortly after they become inactive. As noted previously, 4chan Corpus has toxicity and mysogynist hate comparable to the worst data in Manosphere Corpus and hatespeech above Gab Corpus. We sample this corpus to measure fit to types of discourse and toxicity that can arise from anonymous posting. We concatenate posts in a thread together with post metadata prepended as a header. This source has no marked domains.
E.1 Removed sources
Two additional sources were included in early versions of Paloma, but were removed as access restrictions on these datasets prevent us from rehosting them. We nevertheless present their details here as we still share our findings on these datasets in this Appendix as auxiliary results not part of Paloma.
The Pile
Gao et al. [2020] curate a pretraining corpus from 22 domains in one of the first large open corpora to include mostly non-webscraped text, such as archives of novels or academic papers. It is also explicitly framed as a language modeling benchmark with instructions for standardized evaluations on the validation and test sets, and several open source models have been trained on it [Wang and Komatsuzaki, 2021, Black et al., 2022, Biderman et al., 2023]. It has 22 domains (see Table 15).
Table 15: Domains in The Pile
ICE
Local research teams following guidelines established in Greenbaum and Nelson [1996] collected corpora of English from Canada, East Africa (Kenya & Tanzania), Hong Kong, India, Ireland, Jamaica, Philippines, Singapore, and the USA. Each of these samples of English from around the world is further split into a written and transcribed spoken corpus, except for USA which only has written data (see Table 16). We follow HELM [Liang et al., 2022] in utilizing this corpus to measure disparate performance between these dialects. To permit comparability to HELM, we follow the same preprocessing which leaves in some XML-style tags marking phenomena such as speaker turns.
Table 16: Domains in ICE
Appendix F Reweighting Perplexities
@@
$\text{NLL}_{\text{d,c}}=-\frac{1}{\mathbf{T}(N_{d,c})}\sum_{t\in N_{d,c}}\sum_{i=1}^{\mid t\mid}\text{ln}~{}p(t_{i}|t_{<i})$
@@
$\alpha_{d,c^{\prime}}=\frac{\mathbf{T}(N_{d,c^{\prime}})}{\sum\limits_{d^{\prime}\in D}\mathbf{T}(N_{d^{\prime},c^{\prime}})}$
@@
$\text{perplexity}=\text{exp}\left(\sum_{d\in D}\alpha_{d,c^{\prime}}\text{NLL}_{\text{d,c}}\right)$
Appendix G Baseline Models
The 6 baseline 1B parameter models that we train employ the following architecture: 2048 maximum sequence length, 2048 model dimension, 16 layers, 16 attention heads, RoPE embedding [Su et al., 2021], SwiGLU activation [Shazeer, 2020], mixed precision, non-parametric layer normalization, and sequential model blocks for attention and feed-forward networks. We use EleutherAI’s GPT NeoX tokenizer [Black et al., 2022] but add 3 additional special tokens that are used to mask PII in Dolma. We train to 35k steps ($\sim$150B tokens) with the following LionW optimizer [Chen et al., 2023] configurations: 2.0e-4 peak learning rate, warm-up of 2000 steps, cosine decay to 70k steps ($\sim$300B tokens), 0.1 weight decay, and betas of 0.9 and 0.95. Note that our batch size varies slightly to accommodate two groups of baselines that were run on different hardware. The Dolma and Falcon Refinedweb baselines were run with a batch size of 2112 training instances per step on 24 A100s for 9 days per model. The RedPajama, The Pile, C4, and mC4-en baselines were run with a batch size of 2048 on 64 AMD Instinct MI250X GPUs for 2 days per model. In each case we save model checkpoints every 5k steps ($\sim$20B tokens).
We also include baseline results from the Pythia models [Biderman et al., 2023]. These models do not conform with training guidelines (§3). They do, however, use the GPTNeoX-20B tokenizer [Black et al., 2022] which has an identical vocabulary to our own baseline models, except lacking 3 special tokens used in Dolma. Another similarity is that the Pythia models also have a learning rate schedule set to end at 300B tokens seen, though they train for the full 300B tokens while we train for just 150B tokens of that schedule. This permits comparison between partially trained checkpoints.
Appendix H Formatting and Subsampling
@@
We find preliminary evidence that the monotonic decrease in variability with increased evaluation or training data (see Appendix C.2.1) depends on using the non-concatenated inference input format detailed in Appendix C.2.3. In Table 17 we see that the previously observed trends break down when inputs are concatenated. Additionally, the concatenated documents are drawn from 4 random shufflings where the 4 seeds change for each cell. For comparison the bottom of the table shows results when documents are evaluated separately and with the same set of 4 random seeds for all cells. In both input formats documents that are longer than the model context window are split into separate inputs with no overlap.
@@
Appendix I Most and Least Improved Domains
In Appendix D.1 we show that improvement of LM fit when scaling is unequal from domain to domain. Differences in improvement rates can actually indicate several different training dynamics, exemplified in Figure 7. Looking at performance curves over the underlying factor of scale, helps show more specifically what is going on. Examining the domains at the extreme values of improvement rate is one way to surface interesting details of model fit. In Figure 13 we examine performance curves of the most and least improved domains with respect to number of tokens seen, $\Delta_{t}(\sim 20B,\sim 150B)$, and in Figure 14 we examine the most and least improved with respect to number of model parameters, $\Delta_{p}(85M,805M)$ and $\Delta_{p}(805M,6.4B)$.
Figure 13: Perplexity curves for the most and least improved domains over an increase in tokens seen (See Appendix D.1.1). Columns are specific baseline models; rows are specific evaluation sources.
Figure 14: Perplexity curves for the most and least improved domains over an increase in model size (See Appendix D.1.2). Columns are comparisons of specific model sizes. Each row shows first one (left two subplots) and then another (right two subplots) set of evaluation sources.